A bullet fired from a point on horizontal ground at an angle 30 degree with the horizontal hits the ground 3 km away. By adjusting the angle of projectio, can one hope to hit a target 5 km away?

Grade:11

1 Answers

Piyush Kumar Behera
7 years ago
Solution:
Given that:
Range, R = 3 km
Angle of projection, θ = 30°
Acceleration due to gravity, g = 10 m/s
Horizontal range for the projection velocity u , is given by the relation:
u2/g=2\sqrt{3 }
The maximum range (R ) is achieved by the bullet when it is fired at an angle of 45° with the horizontal, that is,
Rmax=u2/g
On comparing equations (i) and (ii), we get:
Rmax=3\sqrt{3 }=3.46km
Hence, the bullet will not hit a target 5 km away